OpenStudy (anonymous):

Given that f(x) = 3x + 1 and g(x) = 4x+2/3 , solve for g(f(0)). 2 6 8 9

OpenStudy (anonymous):

If you plug 0 in for f(x) you get 1 for that equation then you plug 1 into the other equation giving you 4(1)+2/3=4+2/3=6/3=2

OpenStudy (anonymous):

\[g(f(x)) = 4(3x+1)+\frac{ 2 }{ 3 }\] Now find g(f(0)) by plugging 0 into the equation.

OpenStudy (anonymous):

its answer is 14/3

OpenStudy (anonymous):

g(x) = 4(1)+2/3=4+2/3=12+2/3=14/3

OpenStudy (anonymous):

The best way to solve this function is: \[g[f(x)]=4*f(x)+\frac{2}{3}\]\[f(x)=3x+1\]\[f(0)=1\]So\[g[f(0)]=g(1)\]\[g(1)=4*1+\frac{2}{3}\]\[\boxed{g(x)=\frac{14}{3}}\]

OpenStudy (anonymous):

Well it could be divided by the whole thing but she hasn't specified with any parenthesis or anything so we can only assume.

OpenStudy (anonymous):

its 4x+2 over 3

OpenStudy (anonymous):

OMG that's right... I guess \[g(x)=\frac{4x+2}{3}\]In this case the answer is \[g(1)=\frac{4*1+2}{3}\]\[\boxed{g[f(0)]=2}\]

OpenStudy (anonymous):

They meant the f(x) part

OpenStudy (anonymous):

There we go @D3xt3R :)

OpenStudy (anonymous):

Ya guys can't be overcomplicating things :P

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

haha wow joshes first comment tho